Calculate the calorific value of coal sample from the following data:

Mass of coal: 1g, water and water equivalent of calorimeter: 2 Kg

Specific heat of water: 4.187 KJKg-1C

-1

and rise in temperature: 4.8oC


1
Expert's answer
2021-06-22T07:13:40-0400

= 4.187 × 1 × 2000

= 8374/change in temperature

= 8374/4.8

= 1744.583/1000

= 1.745 Kg/Joules